hello i am doing basic algerbra equations and i am working with word problems and the catorgire is solving equations with variables on both sides of the equation and if possible could you do the eqation in order step by step so i will understand how to do it and i am new at this


If you work christmas day,your company will pay you double time plus a $50 dollar bonus. If you earn $218 for an 8 hour day, what is your hourly rate of pay?

Answer by rfer(16322)

218-50=168/8=$21per hr